OpenStudy (anonymous):

find the slope of each line y-1=0

OpenStudy (anonymous):

This is a horizontal: what is its rise/run?

OpenStudy (anonymous):

what u mean

OpenStudy (anonymous):

Slope is rise over run, right? So how many units does a horizontal line RISE per unit of the domain (don't worry about the size of this unit: the answer will turn out to be very simple)

OpenStudy (anonymous):

i dont understand can u guide me through the steps please

OpenStudy (anonymous):

I'm having some technical difficulties on my end, so, if you haven't already, draw a horizontal line. The definition of slope on a line with two points given is the difference of the y values over the difference of the x values. Just concentrate on the y values: do they ever change, or, since the equation can be rewritten as y=1, is the numerator of your slope equation going to be 1-1 or 0?

OpenStudy (anonymous):

solve for y to get it in standard form y=mx+b

OpenStudy (anonymous):

so do i subtract the y

OpenStudy (anonymous):

well you want to solve for y so you can subtract the y from both sides and get -y=-1, then you would divide the negative or you can add 1 to both sides and get y=1

OpenStudy (anonymous):

it's all the same I guess.

OpenStudy (anonymous):

and that would be my answer

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you're looking to make that equation into a familiar form.

OpenStudy (anonymous):

for example if I had 2y-4x=2 I would want to solve for y to get y=2x+1

OpenStudy (anonymous):

that's just like y=mx+b

OpenStudy (anonymous):

so m=2 x=x and b=1

OpenStudy (anonymous):

so once you get y=1, what do you notice?

OpenStudy (anonymous):

let's pretend it's y=mx+1, what would m have to be to get y=1?

OpenStudy (anonymous):

y=1x+0

OpenStudy (anonymous):

not quite.

OpenStudy (anonymous):

if y=1 and the standard form is y=mx+1 then mx=0, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, now what does m have to equal to get that answer?

OpenStudy (anonymous):

what times x would give you 0?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1*x=0?

OpenStudy (anonymous):

0 times 0

OpenStudy (anonymous):

0*x=0

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

so m would equal to 0

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so the slope is what?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so how should i write it out

OpenStudy (anonymous):

well the slope of y-1=0 is 0

OpenStudy (anonymous):

so zero would be my slope

OpenStudy (anonymous):

yes. :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem.